The Sterling Tire Company's income statement for 20X1 is as follows:


 STERLING TIRE COMPANY Income Statement For the Year Ended December 31, 20X1 


Sales (23,000 tires at $66 each) 

$1,518,000

Variable costs (23,000 tires at $33) 

759,000

Fixed costs 

430,000

Earnings before interest and taxes (EBIT) 

$ 329,000

Interest expense 

51,500

Earnings before taxes (EBT) 

 $277,500
Income tax expense (30%) 83,250
Earnings after taxes (EAT) $ 194,250


a. Compute the degree of operating leverage.

b. Compute the degree of financial leverage.

c. Compute the degree of combined leverage.

d. Compute the break-even point in units.
Answer #1
Answer #2

SOLUTION :


a.


Degree of operating leverage (DOL)

= (Sales - Variable costs) / EBIT

= (1518000 - 759000) / 329000

= 759000 / 329000

= 2.30699

= 2.31  (ANSWER)


b.


Degree of financial leverage (DFL)

= EBIT / EBT 

= 329000 / 277500

= 1.1856

= 1.19(ANSWER).


c.


Degree of combined leverage 

= Contribution margin / EBT  

=  (1518000 - 759000) / 277500

= 2.7351

= 2.74 (ANSWER).


d.


Break-even Point  in units = Fixed costs / Unit contribution margin

= 430000 / (66 - 33)

= 13030.30

= 13030 tiress (ANSWER).


answered by: Tulsiram Garg